OpenStudy (anonymous):

Find approximate solutions of the equation: 2x^2-7x-8=0

OpenStudy (anonymous):

So, a=2 b=-7 c=-8??

OpenStudy (anonymous):

Quadratic formula

OpenStudy (anonymous):

\[x = \frac{7 + \sqrt{113}}{4} , \frac{7 - \sqrt{113}}{4} \]

OpenStudy (anonymous):

So I end up with? -(-7)\[\pm \sqrt{(-7)}^{2}-4*2*-8/2*2\]

OpenStudy (anonymous):

\[7\pm \sqrt{49}+64/4\]

OpenStudy (anonymous):

OK, OK...got it. TY

OpenStudy (anonymous):

So before it said solve and wanted this form. Now its asking for approxiamte solutions. Is that something I can calculate out?

OpenStudy (amistre64):

how close of an approx you need?

OpenStudy (amistre64):

3 is an approx of pi 3.1 is better 3.1415 is even closer 3.14159 even better all are approx.s

OpenStudy (anonymous):

Well it came up with -.91,4.41 I came up with 9.66,434 I must be punching it in the calculator wrong :( But thank you for the definition of approximate

OpenStudy (amistre64):

2x^2-7x-8=0 2x^2-7x = 8 x^2-7/2x = 4 x^2-(7/2)x + (7/4)^2 = 4 + (49/16) (x-(7/4)^2 = 4 + (49/16) x-(7/4) = sqrt(4 + (49/16)) x= (7/4) + sqrt(4 + (49/16)) = 4.407536 = (7/4) - sqrt(4 + (49/16)) = -0.907536 id go with what it came up with :)